Circular No. 6556

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997Y IN NGC 4675 W.-d. Li, Y.-l. Qiu, Q.-y. Qiao, and J.-y. Hu, Beijing Astronomical

Observatory (BAO), on behalf of the BAO supernova survey, report the discovery of an apparent supernova on unfiltered CCD images of NGC 4675 taken on Feb. 2 and 7 with the BAO 0.60-m reflector. SN 1997Y is located at α = 12^{h}45^{m}31^{s}.40, δ = +54^{o}44′17′′.0 (equinox 2000.0), which is 7′′.5 west and 1′′.6 north of the galaxy’s nucleus. The object brightened from mag ∼ 14.8 on Feb. 2 to 14.4 on Feb. 7. Unfiltered CCD images of the same field taken on Jan. 24 showed no star at the position of SN 1997Y.

COMET C/1995 O1 (HALE-BOPP) D. Lynch and R. Russell, The Aerospace Corporation, report: “The

11.2-μm emission feature of comet C/1995 O1 is increasing in strength as the comet approaches the sun. Data taken at the Infrared Telescope Facility and at Mt. Lemmon with the Aerospace BASS spectrograph in 1996 July, Oct. and Nov. reveal that the feature’s contrast above the local 10–11-μm silicate continuum increased from a few percent in July (r = 3.7 AU) to ∼ 15 percent in Nov. (r = 2.25 AU). We also report the detection of a broad (0.2-μm), weak (few percent) emission feature at 11.8 μm. Observers are urged to concentrate on obtaining high S/N-ratio spectra of the comet in the region 8–13 μm in order to monitor the evolution of the features with heliocentric distance.”

GRO J2058+42 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association, on behalf on the RXTE ASM team at GSFC and the Massachusetts Institute of Technology (MIT); A. Peele, GSFC and National Research Council; and R. Remillard, MIT and RXTE ASM team, report: “Analysis of the RXTE ASM light curve of the 195.6-s x-ray pulsar GRO J2058+42 obtained between 1996 Jan. 5 and 1997 Jan. 30 shows the presence of a modulation with a period of ∼ 54 days. This period is half the 110-day period reported from BATSE observations (IAUC 6514), although the ASM is sensitive to lower-energy x-rays in the range 2–12 keV. A sinewave fit to the light curve yields a period of 53.9 ± 1.2 days with epoch of maximum flux at 1996 Oct. 2.4 UT ± 2.0 days. The mean x-ray flux during this period was ∼ 3 mCrab with outbursts reaching ∼ 8 mCrab.”

1997 February 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green